OpenStudy (anonymous):

A positive integer is subtracted from its square. The units digit of the result could be which of the following? a. 3 b. 4 c. 5 d. 6 e. 7

ganeshie8 (ganeshie8):

i would start with testing some positive integers and look for a pattern

ganeshie8 (ganeshie8):

1^2-1 = 0 2^2-2 = 2 3^2-3 = 6 4^2-4 = 12 5^2-5 = 20 6^2-6 = 30 7^2-7 = 42 8^2-8 = 56 .... units digit is always 0, 2, or 6 since 0 and 2 are not there in your options, go with 6

OpenStudy (anonymous):

whats units digit? :[

ganeshie8 (ganeshie8):

ofcourse 6, are you preparing for SAT ?

OpenStudy (anonymous):

yes :[

ganeshie8 (ganeshie8):

okie good luck :)

OpenStudy (anonymous):

thx :]

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:]